# Detectability of Lensed Gravitational Waves in Matched-Filtering Searches

### v1

@@ VI.2 The Role of the Signal Strength
Fig. 5 compares recovered $\mathcal{M}_{c}$ and $\chi_{\rm{eff}}$ distributions for unlensed injections at 125 Mpc (blue) and lensed injections at 125 Mpc (yellow), 625 Mpc (purple), and 1250 Mpc (black). All lensed injection sets exhibit significant deviations from the unlensed distribution at all distances, demonstrating that substantial lensing distortion impacts the accuracy of the recovered source parameters over background noise. However, the discrepancy between lensed and unlensed distributions decreases with increasing source distance, suggesting that matched-filtering searches are less sensitive to large waveform discrepancies in noisier backgrounds. Consequently, we anticipate a smaller SNR loss than observed in the first injection campaign when the signal SNR is lower.
Fig. 6 compares the detectability of unlensed injections (upper) and lensed injections with $y=0.01$ and $M_{\rm{Lz}}=10^{5}M_{\odot}$ (bottom) at distances 125 Mpc (left), 625 Mpc (middle), and 1250 Mpc (right). For the unlensed injections, larger source distances lead to smaller SNR and higher $\xi^{2}$ value. Consequently, 17020, 12878, and 6217 unlensed injections were found for $d_{L}=\{125,625,1250\}$ Mpc. However, lensed injections show an opposite trend with increasing distance. As predicted in our accuracy of recovered source parameters, more lensed injections are recovered with larger matched-filtering SNR compared to the unlensed scenario when the signal strength decreases. As a result, the lensing distortion effect to SNR loss is mitigated in a more noisy background. Moreover, the $\xi^{2}$ value dominates at all distances, indicating that signal strength does not play an important role. High FAR primarily results from large $\xi^{2}$ values due to waveform distortion, regardless of signal strength. Therefore, 1, 8, and 13 lensed injections were found with for $d_{L}=\{125,625,1250\}$ Mpc. This injection campaign shows that matched-filtering rejects signals that significantly deviate from the template more readily at higher SNRs, potentially decreasing detectability even with improved detector sensitivity.
@@ VI.3 General Picture of Detectability of Lensed Gravitational Waves in Matched-Filtering Searches
To illustrate the general impact of lensing on detectability, Fig. 7 displays the matched-filter SNR ratio of lensed to unlensed injections, $\xi^{2}$ value, and FAR in the $y$ vs $M_{\rm{Lz}}$ lens parameter space.
@@
Contrary to optimal SNR analysis predictions that $\rho_{\rm{L}}/\rho_{\rm{UL}}>1$, the top upper panel of Fig. 7 reveals a purple region, where $\rho_{\rm{L}}/\rho_{\rm{UL}}<1$. As shown in the first injection campaign, there is a competing effect between the inaccruacy of the recovered source parameters due to waveform distortion (reduce SNR) and lensing amplification (enhance SNR). Therefore, this purple area corresponds reveal a region where the inaccruacy of the recovered source parameters dominates over the lensing amplification. Note that this region aligned with the $\mathcal{M}\lesssim 0.75$ on the left panel of Fig. 1, further suggesting a possible correlation between matches and the accuracy of the recovered source parameters, which will be further stuided in the future. Outside the purple region, the SNR ratio of lensed to unlensed injections aligns quanlitatively with optimal SNR analysis in the right panel of Fig. 1, indicates the area where lensing amplification dominates the inaccuracy of the recovered source parameters. In addition, the $\rho_{\rm{L}}/\rho_{\rm{UL}}<1$ region shrinks with larger source distances when we compare the SNR ratio from top left to top right panel. It aligned with the result of the second injection campaign, where the impact of lensing distortion in high SNR regime ($\rho_{\rm{opt}}\sim 100$) is the main contribution to the inaccuracy of the recovered source parameters. This contributon from lensing distortion is mitigated by background noises when the signal strength is not prominent ($\rho_{\rm{opt}}\sim 10$).
@@
The middle left panel of Fig. 7 illustrates how different lensing parameters influence the $\xi^{2}$ values at $125$ Mpc. We observe that region with substantial lensing distortion leads to high $\xi^{2}$, which also aligned with the region with $\rho_{\rm{L}}/\rho_{\rm{UL}}<1$ in the top panel in the same figure. This suggest a region of extensive increase in FAR in similar region in the lens parameter space. Furthermore, we observe an expected increase in $\xi^{2}$ at larger source distance due to enhancement in background noises, indicates how background noises can mitigate the impact of waveform distortion in matched-filtering searches.
@@
The bottom left panel of Fig. 7 confirms our expected region with large FAR. aligning with SNR loss and high $\xi^{2}$ regions in the upper and middle panels. This shows that the SNR loss and enhanced $\xi^{2}$ due to strong lensing distortion collectively contribute to the low detectability of lensed gravitational waves. From bottom left to bottom right, there is a overall enhancement in the number of lensed injections with FAR larger than the detection threshold. This is a direct consequence from the impact of weaker signal strength, for which a stronger background noises result in a less significant ranking statistics when SNR, $\xi^{2}$, and all other parameters in Eq. 9 are considered. In addition, we observe that the high FAR region reduces with higher source distances from bottom left to bottom right panels, aligned well with the results in the upper and middle panels.
@@
The third injection campaign yields crucial insights. First, optimal SNR analysis fails to account for the substantial SNR loss due to lensing distortion and incoporate signal consistency tests. These information are critical to access the detectability of lensed gravitational waves in wave optics regime. Lensing analyses and lensing rate calculations, which are based on the optimal SNR method, overestimates the detectability in wave optics regime. Moreover, despite the prominent drop in detectability, lensed signals that significantly deviate from the unlensed template due to wave optics effects are more detectable with weaker signal strength in matched-filtering searches. Consequently, improving detector sensitivity may not enhance the detectability of lensed gravitational waves. Fig. 7 also indicates the need for a lensed gravitational wave template bank, providing valuable guidance on optimal lensing parameter space regions for future template development, particularly for the point mass lens model.
@@ VII Conclusion
A reduction of the detectabilitycan be prominent when the missing physics in the original template bank introduces significant waveform distortions to the gravitational-wave signal, such as beating patterns in gravitational lensing in the wave-dominated regime Takahashi and Nakamura (2003). One conventional approach to approximate the detectability of such signals is to compute the optimal signal-to-noise ratio (SNR). However, this statistic does not account for the possibility of the noise to mimick a real signal. Other detection statistics, such as the signal-consistency test value, are precisley included in the search pipelines to mitigate these effects. A comprehensive study of the background noise in the detectors is also necessary to obtain the probability that a given signal is real. For the first time, we study the detectability of lensed signals with a state-of-the-art matched-filtering search pipeline, capturing the detailed information in the detection statistics and real background noise effects. We “inject” simulated lensed gravitational wave events into data with real noises and retrieve them using a time-domain matched-filtering search pipeline, GstLAL Messick et al. (2017); Sachdev et al. (2019); Cannon et al. (2021); Tsukada et al. (2023).
@@
We perform three distinct injection campaigns to determine the detectability of lensed gravitational waves. The first two injection campaigns study respectively the role of lensing and the role of the signal strength, while the third one investigates the general picture of the detectability of lensed gravitational waves. We observed that matched-filtering searches using the unlensed templates could retrieve a lensed and magnified signal with a higher SNR in the first two injection campaigns, which, for a fixed impact parameter, encompases different lens masses and source distances. However, a fraction of lensed signals lose SNR due to inaccurate recovered source parameters in the search pipeline. This loss in SNR scales with the lensing amplification and the signal strength. One representative scenario is a significant fraction of highly magnified and distorted lensed gravitational waves ($y=0.01$, $M_{\rm{Lz}}=10^{5}M_{\odot}$) at $125$ Mpc (with an unlensed SNR $\sim 100$) is being retrieved with matched-filter SNR $<10$. In addition, the autocorrelation based signal-consistency test value $\xi^{2}$ plays a crucial role. Searching for lensed gravitational waves with an unlensed template bank can increase the $\xi^{2}$ value to dominate the detectability due to substantial waveform distortion. With the lost in SNR and increase in the $\xi^{2}$ value, which can increase the false alarm probability by many orders of magnitude, the detection efficiency of a signal at 125 Mpc drops from $\sim 90\%$ (unlensed) to $<1\%$ (with $y=0.01$, $M_{\rm{Lz}}=10^{5}M_{\odot}$). These two effects, which optimal SNR analyses do not predict, suggest that lensed gravitational waves are less detectable in matched-filtering searches. As shown in the third injection campaign in Sec. VI.3, lensed gravitational waves in the wave optics regime, which introduces the most significant waveform deviation, deviate from optimal SNR analysis. This highlighted that the wave optics regime has the most significant impact on gravitational-wave signals, guiding future efforts in lensed template bank construction.
@@
Based on the insights gained from our injection campaigns, we suggest the following strategies for detecting the lensed gravitational waves. First, matched-filtering searches with a lensed template bank are necessary to probe the lens signatures, especially in the regions of large waveform distortions. Construction of such a template bank requires a new injection campaign that further samples all source parameters, as well as a critical assesment of the computational efficiency of extending the parameter space of the template bank. Second, we suggest that coherent wave burst pipelines can offer an alternative path in detecting lensed gravitational-wave signals due to its model-independent nature Drago et al. (2021). Unlike matched-filtering searches, coherent wave burst can identify loud signals without relying on pre-defined waveform templates, making it more likely to capture lensing-induced modifications. Performing follow-up lensing analyses on gravitational wave candidates found only by coherent wave burst pipeline presents an avenue for detecting lensing signatures that may have been overlooked by template-based searches. This approach complements matched-filtering techniques and could provide a more comprehensive view of lensed gravitational wave events.
@@
Since the detectability of lensed gravitational waves in matched-filtering searches differs from optimal SNR analyses, there are several consequences for other lensing analyses. First, and foremost, the detectability of lensed gravitational waves in the wave-dominated regime is significantly lower in the unlensed matched-filtering searches compared to the optimal SNR approximation. Any follow-up parameter estimation that probes the wave optics signatures can suffer from significant bias, as they are only looking at signals already in current catalogs, and these will not include highly distorted lensed signals. Properly accounting for these selection bias requires a large-scale injection campaign that further samples all the source and lens parameters to specify the probability of detection in the parameter estimation. Besides, reassessing the lensing detection forecasts and constraints based on optimal SNR are necessary. Current lensing rate estimations are based on optimal SNR detection thresholds Meena et al. (2022); Mishra et al. (2021); Meena et al. (2022); Mishra et al. (2024). Although significant lensing distortions are rare, current analyses overestimate the detection rate of lensed gravitational waves. Furthermore, the same applies when setting constraints on the source and lens populations from the (non-)observation of lensing in the gravitational-wave catalogs. This is particularly relevant for compact dark-matter scenarios, which could constitute an effective population of lenses for gravitational waves in the wave-dominated regime. In the past, constraints on those scenarios have been performed based on optimal SNR analyses Jung and Shin (2019); Diego (2020); Liao et al. (2020); Oguri and Takahashi (2020); Urrutia and Vaskonen (2021); Tambalo et al. (2023); Zhou et al. (2022). Since these constraints do not correctly include the reduced detectability in matched-filtering searches shown in our work, the current constraints are overoptimistic in some parts of the parameter space. A full injection campaign is required for these re-analysis.
@@
Searches of new physics in gravitational-wave observations may require dedicated pipelines. This work showcases a scenario in which traditional optimal SNR analysis deviates significantly from matched-filter searches when the new physics are not included in the template. Focusing on the case of gravitational wave lensing, we also show how missing physics impacts the detectability of recovered source parameters and signal consistency tests. Our methodology is however more general and can be applied to other problems of interest. In the future, we will explore this further, performing other injection campaigns to follow up on how different physics, such as deviation from general relativity, impact the detectability. Such investigation can unveil the complex dynamics of how noises and different physics correlate in matched-filtering searches and, hence, help improve the corresponding detection method.

### v2

@@ VI.2 The role of the signal strength
Figure 5 compares the recovered $\mathcal{M}_{c}$ and $\chi_{\rm{eff}}$ distributions for unlensed injections at 125 Mpc (blue) and lensed injections at 125 (purple), 625 (orange), and 1250 Mpc (green). All lensed injection sets exhibit significant deviations from the unlensed distribution at all distances, demonstrating that substantial lensing distortion impacts the accuracy of the recovered source parameters against background noise. However, the discrepancy between lensed and unlensed distributions decreases with increasing source distance, suggesting that matched-filtering searches are less sensitive to large waveform discrepancies in noisier backgrounds. Consequently, we anticipate a smaller SNR loss than observed in the first injection campaign when the signal SNR is lower.
Figure 6 compares the detectability of unlensed injections (upper) and lensed injections with $y=0.01$ and $M_{\rm{Lz}}=10^{5}M_{\odot}$ (bottom) at distances of 125 (left), 625 (middle), and 1250 Mpc (right). For the unlensed injections, larger source distances lead to smaller SNR and higher $\xi^{2}$ values. Consequently, 17020, 12878, and 6217 unlensed injections were found for $d_{L}={125,625,1250}$ Mpc. However, lensed injections exhibit an opposite trend with increasing distance. As predicted in our analysis of the accuracy of recovered source parameters, more lensed injections are recovered with larger matched-filtering SNRs compared to the unlensed scenario when the signal strength decreases. As a result, the effect of lensing distortion on SNR loss is mitigated in a noisier background.
Moreover, the $\xi^{2}$ value dominates at all distances, indicating that signal strength does not play an important role. High FAR values primarily result from large $\xi^{2}$ values caused by waveform distortion, regardless of signal strength. Therefore, 1, 8, and 13 lensed injections were found for $d_{L}={125,625,1250}$ Mpc. This injection campaign demonstrates that matched filtering more readily rejects signals that significantly deviate from the template at higher SNRs, potentially decreasing detectability even with improved detector sensitivity.
@@ VI.3 General picture of detectability of lensed gravitational waves in matched-filtering searches
To illustrate the general impact of lensing on detectability, Fig. 7 displays the matched-filter SNR ratio of lensed to unlensed injections, the $\xi^{2}$ value, and the FAR within the $y$ vs. $M_{\rm{Lz}}$ lens parameter space.
@@
Contrary to predictions from optimal SNR analysis that $\rho_{\rm{L}}/\rho_{\rm{UL}}>1$, the top upper panel of Fig. 7 reveals a purple region where $\rho_{\rm{L}}/\rho_{\rm{UL}}<1$. As shown in the first injection campaign, there is a competing effect between the inaccuracy of the recovered source parameters caused by waveform distortion (reducing SNR) and lensing amplification (enhancing SNR). Therefore, this purple area underscores a region where the inaccuracy of the recovered source parameters dominates over the lensing amplification. Notably, this region aligns with $\mathcal{M}\lesssim 0.75$ on the left panel of Fig. 1, further suggesting a possible correlation between matches and the accuracy of the recovered source parameters, which will be explored in future studies.
Outside the purple region, the SNR ratio of lensed to unlensed injections aligns quantitatively with optimal SNR analysis. The right panel of Fig. 1 indicates an area where lensing amplification dominates over the inaccuracy of the recovered source parameters. In addition, the $\rho_{\rm{L}}/\rho_{\rm{UL}}<1$ region shrinks with increasing source distances when we compare the SNR ratios from the top left to the top right panel. This observation aligns with the results of the second injection campaign, where the impact of lensing distortion in the high-SNR regime ($\rho_{\rm{opt}}\sim 100$) is the primary contributor to the inaccuracy of the recovered source parameters. This contribution from lensing distortion is mitigated by background noise when the signal strength is less prominent ($\rho_{\rm{opt}}\sim 10$).
@@
The middle left panel of Fig. 7 illustrates how different lensing parameters influence the $\xi^{2}$ values at $125$ Mpc. We observe that the region with substantial lensing distortion leads to high $\xi^{2}$ values, which also aligns with the region where $\rho_{\rm{L}}/\rho_{\rm{UL}}<1$ in the top panel of the same figure. This suggests a region of extensive increases in FAR within similar regions of the lens parameter space. Furthermore, we observe an expected increase in $\xi^{2}$ at larger source distances due to enhanced background noise, indicating how background noise can mitigate the impact of waveform distortion in matched-filtering searches.
@@
The bottom left panel of Fig. 7 confirms our expected region with a large FAR, aligning with the SNR loss and high $\xi^{2}$ regions in the upper and middle panels. This demonstrates that the SNR loss and enhanced $\xi^{2}$ caused by strong lensing distortion collectively contribute to the low detectability of lensed gravitational waves. From the bottom left to the bottom right panel, there is an overall increase in the number of lensed injections with FAR values larger than the detection threshold. This is a direct consequence of the impact of weaker signal strength, where stronger background noise results in less significant ranking statistics when SNR, $\xi^{2}$, and all other parameters in Eq.( 9) are considered. In addition, we observe that the high FAR region decreases with increasing source distances from the bottom left to the bottom right panels, which aligns well with the results in the upper and middle panels.
@@
The third injection campaign yields crucial insights. First, optimal SNR analysis fails to account for the substantial SNR loss caused by lensing distortion and the incorporation of signal-consistency tests. This information is critical for assessing the detectability of lensed gravitational waves in wave optics regimes. Lensing analyses and lensing rate calculations, which are based on the optimal SNR method, overestimate detectability in the wave optics regime. Moreover, despite the prominent drop in detectability, lensed signals that significantly deviate from the unlensed template due to wave optics effects are more detectable with weaker signal strength in matched-filtering searches. Consequently, improving detector sensitivity may not enhance the detectability of lensed gravitational waves. Figure 7 also highlights the need for a lensed gravitational-wave template bank, offering valuable guidance on optimal lensing parameter space regions for future template development, particularly for the point mass lens model.
@@ VII Conclusion
A reduction of the detectability can be prominent when the missing physics in the original template bank introduces significant waveform distortions to the gravitational-wave signal, such as beating patterns in gravitational lensing in the wave-dominated regime Takahashi and Nakamura (2003). One conventional approach to approximate the detectability of such signals is to compute the optimal SNR. However, this statistic does not account for the possibility of the noise mimicking a real signal. Other detection statistics, such as the signal-consistency test value, are precisely included in the search pipelines to mitigate these effects. A comprehensive study of the background noise in the detectors is also necessary to obtain the probability that a given signal is real.
For the first time, we study the detectability of lensed signals with a state-of-the-art matched-filtering search pipeline, capturing the detailed information in the detection statistics and real background noise effects. We “inject” simulated lensed gravitational-wave events into data with real noises and retrieve them using a time-domain matched-filtering search pipeline, GstLAL Messick et al. (2017); Sachdev et al. (2019); Cannon et al. (2021); Tsukada et al. (2023).
@@
We perform three distinct injection campaigns to determine the detectability of lensed gravitational waves. The first two injection campaigns study, respectively, the role of lensing and the role of the signal strength, while the third one investigates the general picture of the detectability of lensed gravitational waves. We observed that matched-filtering searches using the unlensed templates could retrieve a lensed and magnified signal with a higher SNR in the first two injection campaigns, which, for a fixed impact parameter, encompass different lens masses and source distances. However, a fraction of lensed signals loses SNR due to inaccurate recovered source parameters in the search pipeline. The number of signals with a decreased SNR scales with both the lensing amplification and the signal strength. One representative scenario is a significant fraction of highly magnified and distorted lensed gravitational waves ($y=0.01$, $M_{\rm{Lz}}=10^{5}M_{\odot}$) at $125$ Mpc (with an unlensed SNR $\sim 100$) being retrieved with matched-filter SNR $<10$.
In addition, the autocorrelation-based signal-consistency test value $\xi^{2}$ plays a crucial role. Searching for lensed gravitational waves with an unlensed template bank can increase the $\xi^{2}$ value to dominate the detectability due to substantial waveform distortion. With the loss in SNR and increase in the $\xi^{2}$ value, which can increase the false alarm probability by many orders of magnitude, the detection efficiency of a signal at 125 Mpc drops from $\sim 90\%$ (unlensed) to $<1\%$ (with $y=0.01$, $M_{\rm{Lz}}=10^{5}M_{\odot}$). These two effects, which optimal SNR analyses do not predict, suggest that lensed gravitational waves are less detectable in matched-filtering searches.
As shown in the third injection campaign in Sec. VI.3, lensed gravitational waves in the wave optics regime, which introduces the most significant waveform deviation, deviate from optimal SNR analysis. This highlighted that the wave optics regime has the most significant impact on gravitational-wave signals, guiding future efforts in lensed template bank construction.
@@
Based on the insights gained from our injection campaigns, we suggest the following strategies for detecting the lensed gravitational waves. First, matched-filtering searches with a lensed template bank are necessary to probe the lens signatures, especially in the regions of large waveform distortions. Construction of such a template bank requires a new injection campaign that further samples all source parameters, as well as a critical assessment of the computational efficiency of extending the parameter space of the template bank. Second, we suggest that coherent wave burst pipelines can offer an alternative path in detecting lensed gravitational-wave signals due to its model-independent nature Drago et al. (2021). Unlike matched-filtering searches, coherent wave bursts can identify loud signals without relying on predefined waveform templates, making it more likely to capture lensing-induced modifications. Performing follow-up lensing analyses on gravitational-wave candidates found only by coherent wave burst pipeline presents an avenue for detecting lensing signatures that may have been overlooked by template-based searches. This approach complements matched-filtering techniques and could provide a more comprehensive view of lensed gravitational-wave events.
@@
Since the detectability of lensed gravitational waves in matched-filtering searches differs from optimal SNR analyses, there are several consequences for other lensing analyses. First and foremost, the detectability of lensed gravitational waves in the wave-dominated regime is significantly lower in the unlensed matched-filtering searches compared to the optimal SNR approximation. Any follow-up parameter estimation that probes the wave optics signatures can suffer from significant bias, as they are only looking at signals already in current catalogs, and these will not include highly distorted lensed signals. Properly accounting for these selection biases requires a large-scale injection campaign that further samples all the source and lens parameters to specify the probability of detection in the parameter estimation.
Additionally, reassessing the lensing detection forecasts and constraints based on optimal SNR is necessary. Current lensing rate estimations are based on optimal SNR detection thresholds Meena et al. (2022); Mishra et al. (2021); Meena et al. (2022); Mishra et al. (2024). Although significant lensing distortions are rare, current analyses overestimate the detection rate of lensed gravitational waves. Furthermore, the same applies when setting constraints on the source and lens populations from the (non)observation of lensing in the gravitational-wave catalogs. This is particularly relevant for compact dark matter scenarios, which could constitute an effective population of lenses for gravitational waves in the wave-dominated regime. In the past, constraints on those scenarios have been performed based on optimal SNR analyses Jung and Shin (2019); Diego (2020); Liao et al. (2020); Oguri and Takahashi (2020); Urrutia and Vaskonen (2021); Tambalo et al. (2023); Zhou et al. (2022). Since these constraints do not correctly include the reduced detectability in matched-filtering searches shown in our work, the current constraints are overoptimistic in some parts of the parameter space. A full injection campaign is required for these reanalysis.
@@
Searches of new physics in gravitational-wave observations may require dedicated pipelines. This work showcases a scenario in which traditional optimal SNR analysis deviates significantly from matched-filter searches when the new physics are not included in the template. Focusing on the case of gravitational-wave lensing, we also show how missing physics impacts the detectability of recovered source parameters and signal-consistency tests. Our methodology is, however, more general and can be applied to other problems of interest. In the future, we will explore this further, performing other injection campaigns to follow up on how different physics, such as deviation from general relativity, impact the detectability. Such investigation can unveil the complex dynamics of how noises and different physics correlate in matched-filtering searches and, hence, help improve the corresponding detection method.